The room is 2 meters longer than it is wide it's perimeter is 24 meters . Find the width in meters.
Answer by checkley77(12844)

L=W+2
2L+2W=24
2(W+2)+2W=24
2W+4+2W=24
4W=24-4
4W=20
W=20/4
W=5 ANS.
L=5+2=7 ANS. FOR THE LENGTH.
PROOF:
2*7+2*5=24
14+10=24